# HMID-Net: An Exploration of Masked Image Modeling and Knowledge Distillation in Hyperbolic Space

### v1

@@ Abstract
keywords:
Visual-Semantic hierarchy , Hyperbolic Space , Knowledge Distillation , Masked Image Modeling , Multi-modal learning
journal: Pattern Recognition
[…]
[hit]organization=Faculty of Computing, Harbin Institute of Technology,city=Harbin, postcode=150001, country=China
[…]
[hust]organization=College of Computer Science and Technology, Harbin University of Science and Technology, city=Harbin, postcode=150001, country=China
{graphicalabstract}
{highlights}
We propose HMID-Net, the first MIM and knowledge distillation in hyperbolic space.
We introduce Feature Interaction Distillation in hyperbolic space and its loss function.
We are the first to show that MIM and knowledge distillation in hyperbolic space achieve success.
Experiments demonstrate notable improvements across various tasks.
1 Introduction
Humans can perceive the real world through images, where a single image encapsulates a wealth of information. This information can be articulated through diverse textual descriptions, each providing a distinct interpretation. These diverse descriptions exhibit multiple hierarchical relationships. As humans, we possess the ability to reason from each description and organize the information into coherent visual-semantic hierarchy [1, 2].
@@ 1 Introduction
In recent years, the rapid progression of deep learning has been predominantly fueled by substantial advancements in hardware, enabling the feasibility of large-scale pre-trained Vision-Language Models (VLMs). Multimodal Large Language Models (MLLMs) have emerged as a central focus of contemporary research. A range of multimodal models like CLIP [3] and ALIGN [4] have emerged and achieved remarkable success on various downstream tasks, such as detection [5, 6], classification [3], and retrieval [7, 8, 9]. In particular, CLIP is trained on a dataset consisting of approximately 400 million image-text pairs, while ALIGN is trained on 1.8 billion image-text pairs, pioneering a new pretraining paradigm and enabling these models to perform a variety of tasks without the need for fine-tuning. This raises a question: in the absence of substantial data, how can we effectively train a high-performance model? Several researchers have adopted techniques such as knowledge distillation [10, 11, 12], prompt tuning [13, 14], and adapter [15, 16] to mitigate training costs.
However, these methods are all based on Euclidean space, where the capacity of embeddings is linearly tied to their dimensionality, which limits their ability to effectively capture complex data relationships, such as the visual-semantic hierarchy. In contrast, in hyperbolic space, the capacity increases exponentially with the radius of the sphere, enabling it to accommodate embeddings of any structure while preserving their inherent properties [17]. Currently, hyperbolic space has been widely applied in various fields, including classification [18, 19, 20, 21], segmentation [22], detection [23], retrieval [2, 24], and point cloud [25]. Similarly, in hyperbolic space, how can we train a more efficient model under conditions of limited data? While knowledge distillation have proven effective in Euclidean space, their potential remains underexplored in hyperbolic space. We identify the following potential reasons: (1) Most deep learning frameworks and libraries are predominantly designed for Euclidean space, offering limited support for hyperbolic space, significantly increasing the technical difficulty. (2) Due to the inherent differences in distance measurement between Euclidean and hyperbolic spaces, designing a loss function in hyperbolic space presents unique challenges. (3) The fundamental operations in hyperbolic space are more intricate and demand significantly greater computational resources.
@@
2 Realted Works
2.1 Masked Image Modeling
@@
Given the remarkable success of the Masked Language Model (MLM) in Natural Language Processing (NLP), researchers have extensively explored and investigated analogous approaches in vision [26, 27, 28, 29]. MAE [26] constructs an asymmetric encoder-decoder framework, where the encoder randomly masks and shuffles 75% of the image, and the decoder is responsible for reconstructing the original pixels. [30] demonstrates through theoretical derivation that MAE can implicitly align masked and unmasked views. FLIP [31] applies MAE to multimodal learning and demonstrates that the reconstruction loss and text masking are not necessary.
2.2 Prompt tuning, Adapter and Knowledge Distillation
@@
Prompt tuning is a text input segment, such as “a photo of the large {}”, that guides a pretrained language model to generate specific outputs or perform tasks. It enables task-solving without traditional fine-tuning. However, crafting effective manual prompts requires expertise and is highly time-consuming. CoOp [13] proposes two learnable prompts: Unified Context and Class-Specific Context, outperforming manual prompts across various domains. However, CoOp faces challenges with generalization to unseen categories, a limitation that CoCoOp [14] attributes to overfitting.
@@
Adapter is a plug-and-play neural network module that requires training only small additional components. Tip-Adapter [15] utilizes a query-key caching mechanism, eliminating the need for additional training. CLIP-Adapter [32] proposes integrating an adapter at the end of the backbone network, instead of utilizing prompts, enabling few-shot fine-tuning of the model. CALIP [33] enhances CLIP with a parameter-free attention module for cross-modal interaction, eliminating the need for additional downstream data or training.
@@
Knowledge Distillation transfers generalization features from a teacher model to a student model, enabling high performance with reduced computational cost. [34] introduces self distillation, where the model serves as both the teacher and the student. TinyCLIP [10] introduces affinity imitation and weight inheritance, effectively reducing model size and applying knowledge distillation to CLIP for the first time. CLIP-KD [11] validates the effectiveness of CLIP knowledge distillation from the perspectives of relationships, features, gradients, and contrastive modes. However, these methods have not been explored in hyperbolic space. In this paper, we investigate knowledge distillation within hyperbolic space.
2.3 Hyperbolic deep neural networks
@@
In hyperbolic space, there are five well-known isometric models: the Lorentz (Hyperboloid) model, the Poincaré ball model, the Poincaré half-space model, the Klein model, and the Hemisphere model [35]. [36] uses the Poincaré ball to model hierarchical relationships and applies Riemannian gradient optimization for training. [37] reconstructs Euclidean operations (addition, multiplication, FFN) in hyperbolic space. [17] introduces entailment cones to establish a partial order and express entailment in the Poincaré ball model. In computer vision, hyperbolic space has a wide range of applications [18, 19, 20, 21, 23, 22]. MERU [2] is the first to integrate hyperbolic space into vision-language models (VLMs), with the aim of capturing the visual-semantic hierarchy depicted in Fig. 1(a). However, the aforementioned methods cannot directly leverage pre-trained models in Euclidean space. In this paper, we employ MIM and knowledge distillation to train an efficient model in hyperbolic space.
3 Preliminary
@@
Hyperbolic geometry is a special case of Riemannian geometry. Before presenting our method, this section first introduces Riemannian geometry (Section 3.1) and Lorentz model (Section 3.2).
3.1 Riemannian geometry
@@ 3.2 Lorentz model
4 Approach
4.1 A Review of CLIP
CLIP [3] leverages contrastive learning to project images and text into a shared semantic space, allowing the model to effectively capture and understand their semantic relationships. Specifically, CLIP employs a dual-tower architecture, processing image and text independently through two encoders: the image encoder $h_{\theta}$ and the text encoder $g_{\theta}$. For an image $I\in\mathbb{R}^{H\times W\times 3}$, the image encoder utilizes either ResNet [38] or ViT [39] to extract image feature, denoted as $f(v)=h_{\theta}(I)$. Similarly, for a text $T$, the text encoder utilizes a Transformer model to convert the text into text features, denoted as $f(l)=g_{\theta}(T)$. During training, image-text pairs within the same batch are positive samples, while unpaired images and text are negative samples. The model uses a contrastive loss to maximize the alignment between images and their corresponding textual descriptions in the shared semantic space. The similarity between the image and text embeddings is computed, typically using cosine similarity, and the contrastive loss function $\mathcal{L}_{cl}$ is defined as Eq. 7:
$sim(f(v),f(l))=\frac{f(v)\cdot f(l)}{\|f(v)\|\|f(l)\|}$ (6)
@@ 4.1 A Review of CLIP
4.2 Overview of our proposed method
Fig. 2(a) depicts the training process of our method. The architecture consists of a student and a teacher model. For a given image-text pair $(I,T)$, the teacher and student models generate embeddings $[f_{\textbf{T}}(v),f_{\textbf{T}}(l)]$ and $[f_{\textbf{S}}(v),f_{\textbf{S}}(l)]$, respectively. These embeddings are then projected into hyperbolic space as $[f^{{}^{\prime}}_{\textbf{T}}(v),f^{{}^{\prime}}_{\textbf{T}}(l)]$ and $[f^{{}^{\prime}}_{\textbf{S}}(v),f^{{}^{\prime}}_{\textbf{S}}(l)]$. In hyperbolic space, contrastive learning is applied between $f^{{}^{\prime}}_{\textbf{S}}(v)$ and $f^{{}^{\prime}}_{\textbf{S}}(l)$, while the contrastive loss aligns the image-text pairs and the entailment loss forces the model to learn the visual-semantic hierarchy, as detailed in Sections 4.4 and 4.5.
@@ 4.2 Overview of our proposed method
4.3 Masked image
We adopt Vision Transformer (ViT) [39] as the image encoder and the Transformer as the text encoder. For a given image, it is initially partitioned into non-overlapping patches, with a substantial portion (e.g., 50%) of the patches randomly masked. Only the unmasked patches are input into the network, following [26, 31]. In this paper, we do not reconstruct the original pixels, as noted in [31], because it has minimal impact on the final results and introduces unnecessary computational complexity.
4.4 Hyperbolic contrastive learning
@@
4.5 Feature Interaction Distillation
To explore effective knowledge distillation methods, [11] first introduced interactive contrastive learning. In this paper, we extend this method to the hyperbolic space. Specifically, in the hyperbolic space, given the student image embedding $f_{\textbf{S}}^{{}^{\prime}}(v)$, student text embedding $f_{\textbf{S}}^{{}^{\prime}}(l)$, teacher text embedding $f_{\textbf{T}}^{{}^{\prime}}(l)$, and teacher image embedding $f_{\textbf{T}}^{{}^{\prime}}(v)$, we replace $f_{\textbf{S}}^{{}^{\prime}}(l)$ with $f_{\textbf{T}}^{{}^{\prime}}(l)$ when calculating the image-to-text (I2T) contrastive learning loss, which can be formulated as:
@@ 4.5 Feature Interaction Distillation
4.6 Entailment loss
[1] proposes utilizing partial orders to represent the relationship between vision and text. [17] introduces the Entailment loss to learn image-text pairs and their partial order relationships. Similar to [2], we incorporate the entailment loss to enhance the model’s ability to capture visual-semantic hierarchy.
@@ 4.7 Overall Loss
5 Experiments
In this section, we assess the performance of our method across diverse downstream vision-language (V+L) tasks. Additionally, in Section 5.4, we conduct comprehensive ablation studies to assess the impact of each component on the overall performance.
Dataset Classes Train Val Task ImageNet [40] 1000 1,281,167 50,000 General object classification Food101 [41] 101 75,750 25,250 Fine-grained classification CIFAR10 [42] 10 50,000 10,000 General object classification CIFAR100 [42] 100 50,000 10,000 General object classification SUN397 [43] 397 76,128 19,849 Scene recognition Aircraft [44] 100 3,334 3,333 Fine-grained classification DTD [45] 47 1,880 1,880 Fine-grained classification Pets [46] 37 3,680 3,669 Fine-grained classification Caltech101 [47] 102 3,060 6,084 General object classification Flowers [48] 102 1,020 6,149 Fine-grained classification STL10 [49] 10 5,000 8,000 General object classification Resisc45 [50] 45   25,200 Remote sensing classification Country211 [3] 211   21,100 General object classification MNIST [51] 10 60,000 10,000 Handwritten digit classification CLEVR [52] 8   5,000 Visual question answering SST2 [3] 2   1,821 Sentiment analysis COCO [53]   118,000 5,000 Image and text retrieval Flickr30K [54]   29,000 1,000 Image and text retrieval
5.1 Implementation details
Baselines. We first compare our method with CLIP [3], which effectively achieves joint multimodal representation by embedding images and text into Euclidean space. The primary focus of our work is a comparison with MERU [2], which explores the representation of images and text in the hyperbolic space for the first time. Similar to MERU, we pretrain our model using the Redcaps dataset, which contains $\sim$12 million image-text pairs collected from 350 manually selected subreddits on Reddit. Due to the absence of some data on the website, we are only able to download $\sim$ 7 million image-text pairs. Consequently, we retrain MERU and CLIP using the available dataset.
Datasets. To evaluate our method, we select 18 datasets, including 16 for image classification and 2 for image-text retrieval. The image classification tasks span a wide range of domains, covering general object classification (ImageNet [40], Caltech101 [47]), fine-grained object classification (Food101 [41], Pets [46], Flowers [48]), scene recognition (Sun397 [43]), and remote sensing classification (Resisc45 [50]), among others. For image-text retrieval tasks, we utilize the COCO [53] and Flickr30K [54] datasets. A comprehensive overview of these datasets is provided in Table 1.
Models. We adopt ViT [39] as the image encoder and select three different variants: ViT-S, ViT-B, and ViT-L. The patch size for all models is set to 16. For the text encoder, we use a 12-layer, 512-dimensional Transformer [55], consistent with the implementation of MERU.
Training details. We utilize the publicly available Open-CLIP [56] as the teacher model. During training, the image and text encoders of the teacher model are frozen, while the parameters of the image and text encoders of the student model are updated. The embeddings are projected into the hyperbolic space using the Exponential map, with both the teacher and student models sharing the same curvature $c$. We use the AdamW [57] optimizer to train the model, with a weight decay of 0.2 and a maximum learning rate of $5\times 10^{-4}$. The learning rate undergoes linear growth during the first 10% of the total iterations, followed by a cosine decay until it reaches zero. All models in this paper are trained for 560,000 iterations (approximately 20 epochs) with a batch size of 256. The implementation is based on PyTorch, and the training is conducted on four NVIDIA GeForce RTX 4090 GPUs.
5.2 Image classification
@@
We assess HMID-Net across 16 image classification benchmarks. Table 2 represents the zero-shot image classification performance of HMID-Net. We report the absolute improvements of our method over the baseline CLIP and MERU, with the backbone being ViT-L/16. The left image in Fig. 4 compares our method with CLIP, while the right image compares it with MERU. It significantly outperforms our baselines, CLIP and MERU, on 13 out of the 16 datasets. Specifically, HMID-Net achieves +7.6% improvement over CLIP on the Caltech101 dataset and +6.7% improvement over MERU on the CIFAR100 dataset. Additionally, HMID-Net achieves an accuracy of 28.6% on the ImageNet dataset, surpassing CLIP and MERU by 5.1% and 4.3%, respectively. On the MNIST and SST2 datasets, HMID-Net achieves performance on par with that of CLIP and MERU. As noted by [2], HMID-Net exhibits relatively suboptimal performance on datasets with fewer covered concepts, such as SST2, which is derived from movie reviews. Pretraining on larger datasets may improve performance. Overall, HMID-Net is highly competitive compared to Euclidean space-based methods.
@@
5.3 Image and text retrieval
5.4 Ablations experiments
Masking ratio. We first conducted a study on image masking ratios based solely on MERU, as shown in Table LABEL:mask. The 0% masking ratio refers to our baseline MERU. A 50% masking ratio yields optimal performance, with a 1.7% improvement on ImageNet and a 1.3% improvement on COCO for the T2I retrieval task. Compared to BERT’s 15% masking ratio, images have significant pixel redundancy, enabling a higher masking ratio. However, a 75% masking ratio results in a performance decline due to the loss of critical information, which hinders contrastive learning. This effect is similar to the findings in FLIP [31]. Unless specified otherwise, we use a default masking ratio of 50%.

### v2

# ViSP: A PPO-Driven Framework for Sarcasm Generation with Contrastive Learning

@@ Abstract
Introduction
Humans can perceive the real world through images, where a single image encapsulates a wealth of information. This information can be articulated through diverse textual descriptions, each providing a distinct interpretation. These diverse descriptions exhibit multiple hierarchical relationships. As humans, we possess the ability to reason from each description and organize the information into coherent visual-semantic hierarchy (Vendrov et al. 2016; Desai et al. 2023).
@@ Introduction
In recent years, the rapid progression of deep learning has been predominantly fueled by substantial advancements in hardware, enabling the feasibility of large-scale pre-trained Vision-Language Models (VLMs). Multimodal Large Language Models (MLLMs) have emerged as a central focus of contemporary research. A range of multimodal models like CLIP (Radford et al. 2021) and ALIGN (Jia et al. 2021) have emerged and achieved remarkable success on various downstream tasks, such as detection (Gu et al. 2022; Li* et al. 2022), classification (Radford et al. 2021), and retrieval (Luo et al. 2022; Zhao et al. 2022; Baldrati et al. 2023). In particular, CLIP is trained on a dataset consisting of approximately 400 million image-text pairs, while ALIGN is trained on 1.8 billion image-text pairs, pioneering a new pretraining paradigm and enabling these models to perform a variety of tasks without the need for fine-tuning. This raises a question: in the absence of substantial data, how can we effectively train a high-performance model? Several researchers have adopted techniques such as knowledge distillation (Wu et al. 2023; Yang et al. 2024; Wu et al. 2025), prompt tuning (Zhou et al. 2022b, a), and adapter (Zhang et al. 2022; Jiang et al. 2025) to mitigate training costs.
However, these methods are all based on Euclidean space, where the capacity of embeddings is linearly tied to their dimensionality, which limits their ability to effectively capture complex data relationships, such as the visual-semantic hierarchy. In contrast, in hyperbolic space, the capacity increases exponentially with the radius of the sphere, enabling it to accommodate embeddings of any structure while preserving their inherent properties (Ganea, Bécigneul, and Hofmann 2018a). Currently, hyperbolic space has been widely applied in various fields, including classification (Khrulkov et al. 2020; Liu et al. 2020; Dhall et al. 2020; Kwon et al. 2024), segmentation (Atigh et al. 2022), detection (Kong et al. 2024), retrieval (Desai et al. 2023; Ramasinghe et al. 2024), and point cloud (Feng et al. 2025). Similarly, in hyperbolic space, how can we train a more efficient model under conditions of limited data? While knowledge distillation have proven effective in Euclidean space, their potential remains underexplored in hyperbolic space. We identify the following potential reasons: (1) Most deep learning frameworks and libraries are predominantly designed for Euclidean space, offering limited support for hyperbolic space, significantly increasing the technical difficulty. (2) Due to the inherent differences in distance measurement between Euclidean and hyperbolic spaces, designing a loss function in hyperbolic space presents unique challenges. (3) The fundamental operations in hyperbolic space are more intricate and demand significantly greater computational resources.
@@
Realted Works
Masked Image Modeling
@@
Given the remarkable success of the Masked Language Model (MLM) in Natural Language Processing (NLP), researchers have extensively explored and investigated analogous approaches in vision (He et al. 2022; Xie et al. 2022; Bao et al. 2022; Wei et al. 2022). MAE (He et al. 2022) constructs an asymmetric encoder-decoder framework, where the encoder randomly masks and shuffles 75% of the image, and the decoder is responsible for reconstructing the original pixels. (Zhang, Wang, and Wang 2022) demonstrates through theoretical derivation that MAE can implicitly align masked and unmasked views. FLIP (Li et al. 2023) applies MAE to multimodal learning and demonstrates that the reconstruction loss and text masking are not necessary.
Prompt tuning, Adapter and Knowledge Distillation
@@
Prompt tuning is a text input segment, such as “a photo of the large {}”, that guides a pretrained language model to generate specific outputs or perform tasks. It enables task-solving without traditional fine-tuning. However, crafting effective manual prompts requires expertise and is highly time-consuming. CoOp (Zhou et al. 2022b) proposes two learnable prompts: Unified Context and Class-Specific Context, outperforming manual prompts across various domains. However, CoOp faces challenges with generalization to unseen categories, a limitation that CoCoOp (Zhou et al. 2022a) attributes to overfitting.
@@
Adapter is a plug-and-play neural network module that requires training only small additional components. Tip-Adapter (Zhang et al. 2022) utilizes a query-key caching mechanism, eliminating the need for additional training. CLIP-Adapter (Gao et al. 2024) proposes integrating an adapter at the end of the backbone network, instead of utilizing prompts, enabling few-shot fine-tuning of the model. CALIP (Guo et al. 2023) enhances CLIP with a parameter-free attention module for cross-modal interaction, eliminating the need for additional downstream data or training.
@@
Knowledge Distillation transfers generalization features from a teacher model to a student model, enabling high performance with reduced computational cost. (Zhang et al. 2019) introduces self distillation, where the model serves as both the teacher and the student. TinyCLIP (Wu et al. 2023) introduces affinity imitation and weight inheritance, effectively reducing model size and applying knowledge distillation to CLIP for the first time. CLIP-KD (Yang et al. 2024) validates the effectiveness of CLIP knowledge distillation from the perspectives of relationships, features, gradients, and contrastive modes. However, these methods have not been explored in hyperbolic space. In this paper, we investigate knowledge distillation within hyperbolic space.
Hyperbolic deep neural networks
@@
In hyperbolic space, there are five well-known isometric models: the Lorentz (Hyperboloid) model, the Poincaré ball model, the Poincaré half-space model, the Klein model, and the Hemisphere model (Peng et al. 2021). (Nickel and Kiela 2017) uses the Poincaré ball to model hierarchical relationships and applies Riemannian gradient optimization for training. (Ganea, Bécigneul, and Hofmann 2018b) reconstructs Euclidean operations (addition, multiplication, FFN) in hyperbolic space. (Ganea, Bécigneul, and Hofmann 2018a) introduces entailment cones to establish a partial order and express entailment in the Poincaré ball model. In computer vision, hyperbolic space has a wide range of applications (Khrulkov et al. 2020; Liu et al. 2020; Dhall et al. 2020; Kwon et al. 2024; Kong et al. 2024; Atigh et al. 2022). MERU (Desai et al. 2023) is the first to integrate hyperbolic space into vision-language models (VLMs), with the aim of capturing the visual-semantic hierarchy depicted in Fig. 1(a). However, the aforementioned methods cannot directly leverage pre-trained models in Euclidean space. In this paper, we employ MIM and knowledge distillation to train an efficient model in hyperbolic space.
Preliminary
@@
Hyperbolic geometry is a special case of Riemannian geometry. Before presenting our method, this section first introduces Riemannian geometry (Section Riemannian geometry) and Lorentz model (Section Lorentz model).
Riemannian geometry
@@
Approach
A Review of CLIP
CLIP (Radford et al. 2021) leverages contrastive learning to project images and text into a shared semantic space, allowing the model to effectively capture and understand their semantic relationships. Specifically, CLIP employs a dual-tower architecture, processing image and text independently through two encoders: the image encoder $h_{\theta}$ and the text encoder $g_{\theta}$. For an image $I\in\mathbb{R}^{H\times W\times 3}$, the image encoder utilizes either ResNet (He et al. 2016) or ViT (Dosovitskiy et al. 2021) to extract image feature, denoted as $f(v)=h_{\theta}(I)$. Similarly, for a text $T$, the text encoder utilizes a Transformer model to convert the text into text features, denoted as $f(l)=g_{\theta}(T)$. During training, image-text pairs within the same batch are positive samples, while unpaired images and text are negative samples. The model uses a contrastive loss to maximize the alignment between images and their corresponding textual descriptions in the shared semantic space. The similarity between the image and text embeddings is computed, typically using cosine similarity, and the contrastive loss function $\mathcal{L}_{cl}$ is defined as Eq. 7:
$sim(f(v),f(l))=\frac{f(v)\cdot f(l)}{\|f(v)\|\|f(l)\|}$ (6)
Overview of our proposed method
Fig. 2(a) depicts the training process of our method. The architecture consists of a student and a teacher model. For a given image-text pair $(I,T)$, the teacher and student models generate embeddings $[f_{\textbf{T}}(v),f_{\textbf{T}}(l)]$ and $[f_{\textbf{S}}(v),f_{\textbf{S}}(l)]$, respectively. These embeddings are then projected into hyperbolic space as $[f^{{}^{\prime}}_{\textbf{T}}(v),f^{{}^{\prime}}_{\textbf{T}}(l)]$ and $[f^{{}^{\prime}}_{\textbf{S}}(v),f^{{}^{\prime}}_{\textbf{S}}(l)]$. In hyperbolic space, contrastive learning is applied between $f^{{}^{\prime}}_{\textbf{S}}(v)$ and $f^{{}^{\prime}}_{\textbf{S}}(l)$, while the contrastive loss aligns the image-text pairs and the entailment loss forces the model to learn the visual-semantic hierarchy, as detailed in Sections Hyperbolic contrastive learning and Feature Interaction Distillation.
@@ A Review of CLIP
Masked image
We adopt Vision Transformer (ViT) (Dosovitskiy et al. 2021) as the image encoder and the Transformer as the text encoder. For a given image, it is initially partitioned into non-overlapping patches, with a substantial portion (e.g., 50%) of the patches randomly masked. Only the unmasked patches are input into the network, following (He et al. 2022; Li et al. 2023). In this paper, we do not reconstruct the original pixels, as noted in (Li et al. 2023), because it has minimal impact on the final results and introduces unnecessary computational complexity.
Hyperbolic contrastive learning
@@
Feature Interaction Distillation
To explore effective knowledge distillation methods, (Yang et al. 2024) first introduced interactive contrastive learning. In this paper, we extend this method to the hyperbolic space. Specifically, in the hyperbolic space, given the student image embedding $f_{\textbf{S}}^{{}^{\prime}}(v)$, student text embedding $f_{\textbf{S}}^{{}^{\prime}}(l)$, teacher text embedding $f_{\textbf{T}}^{{}^{\prime}}(l)$, and teacher image embedding $f_{\textbf{T}}^{{}^{\prime}}(v)$, we replace $f_{\textbf{S}}^{{}^{\prime}}(l)$ with $f_{\textbf{T}}^{{}^{\prime}}(l)$ when calculating the image-to-text (I2T) contrastive learning loss, which can be formulated as:
@@
Dataset Classes Train Val Task ImageNet (Deng et al. 2009) 1000 1,281,167 50,000 General object classification Food101 (Bossard, Guillaumin, and Van Gool 2014) 101 75,750 25,250 Fine-grained classification CIFAR10 (Krizhevsky, Hinton et al. 2009) 10 50,000 10,000 General object classification CIFAR100 (Krizhevsky, Hinton et al. 2009) 100 50,000 10,000 General object classification SUN397 (Xiao et al. 2010) 397 76,128 19,849 Scene recognition Aircraft (Maji et al. 2013) 100 3,334 3,333 Fine-grained classification DTD (Cimpoi et al. 2014) 47 1,880 1,880 Fine-grained classification Pets (Parkhi et al. 2012) 37 3,680 3,669 Fine-grained classification Caltech101 (Fei-Fei, Fergus, and Perona 2004) 102 3,060 6,084 General object classification Flowers (Nilsback and Zisserman 2008) 102 1,020 6,149 Fine-grained classification STL10 (Coates, Ng, and Lee 2011) 10 5,000 8,000 General object classification Resisc45 (Cheng, Han, and Lu 2017) 45   25,200 Remote sensing classification Country211 (Radford et al. 2021) 211   21,100 General object classification MNIST (LeCun et al. 1998) 10 60,000 10,000 Handwritten digit classification CLEVR (Johnson et al. 2017) 8   5,000 Visual question answering SST2 (Radford et al. 2021) 2   1,821 Sentiment analysis COCO (Chen et al. 2015)   118,000 5,000 Image and text retrieval Flickr30K (Young et al. 2014)   29,000 1,000 Image and text retrieval
Entailment loss
(Vendrov et al. 2016) proposes utilizing partial orders to represent the relationship between vision and text. (Ganea, Bécigneul, and Hofmann 2018a) introduces the Entailment loss to learn image-text pairs and their partial order relationships. Similar to (Desai et al. 2023), we incorporate the entailment loss to enhance the model’s ability to capture visual-semantic hierarchy.
@@
Experiments
In this section, we assess the performance of our method across diverse downstream vision-language (V+L) tasks. Additionally, in Section Ablations experiments, we conduct comprehensive ablation studies to assess the impact of each component on the overall performance.
Implementation details
Baselines. We first compare our method with CLIP (Radford et al. 2021), which effectively achieves joint multimodal representation by embedding images and text into Euclidean space. The primary focus of our work is a comparison with MERU (Desai et al. 2023), which explores the representation of images and text in the hyperbolic space for the first time. Similar to MERU, we pretrain our model using the Redcaps dataset, which contains $\sim$12 million image-text pairs collected from 350 manually selected subreddits on Reddit. Due to the absence of some data on the website, we are only able to download $\sim$ 7 million image-text pairs. Consequently, we retrain MERU and CLIP using the available dataset.
Datasets. To evaluate our method, we select 18 datasets, including 16 for image classification and 2 for image-text retrieval. The image classification tasks span a wide range of domains, covering general object classification (ImageNet (Deng et al. 2009), Caltech101 (Fei-Fei, Fergus, and Perona 2004)), fine-grained object classification (Food101 (Bossard, Guillaumin, and Van Gool 2014), Pets (Parkhi et al. 2012), Flowers (Nilsback and Zisserman 2008)), scene recognition (Sun397 (Xiao et al. 2010)), and remote sensing classification (Resisc45 (Cheng, Han, and Lu 2017)), among others. For image-text retrieval tasks, we utilize the COCO (Chen et al. 2015) and Flickr30K (Young et al. 2014) datasets. A comprehensive overview of these datasets is provided in Table 1.
Models. We adopt ViT (Dosovitskiy et al. 2021) as the image encoder and select three different variants: ViT-S, ViT-B, and ViT-L. The patch size for all models is set to 16. For the text encoder, we use a 12-layer, 512-dimensional Transformer (Vaswani et al. 2017), consistent with the implementation of MERU.
Training details. We utilize the publicly available Open-CLIP (Cherti et al. 2023) as the teacher model. During training, the image and text encoders of the teacher model are frozen, while the parameters of the image and text encoders of the student model are updated. The embeddings are projected into the hyperbolic space using the Exponential map, with both the teacher and student models sharing the same curvature $c$. We use the AdamW (Loshchilov and Hutter 2019) optimizer to train the model, with a weight decay of 0.2 and a maximum learning rate of $5\times 10^{-4}$. The learning rate undergoes linear growth during the first 10% of the total iterations, followed by a cosine decay until it reaches zero. All models in this paper are trained for 560,000 iterations (approximately 20 epochs) with a batch size of 256. The implementation is based on PyTorch, and the training is conducted on four NVIDIA GeForce RTX 4090 GPUs.
Image classification
@@
We assess HMID-Net across 16 image classification benchmarks. Table 2 represents the zero-shot image classification performance of HMID-Net. We report the absolute improvements of our method over the baseline CLIP and MERU, with the backbone being ViT-L/16. The left image in Fig. 4 compares our method with CLIP, while the right image compares it with MERU. It significantly outperforms our baselines, CLIP and MERU, on 13 out of the 16 datasets. Specifically, HMID-Net achieves +7.6% improvement over CLIP on the Caltech101 dataset and +6.7% improvement over MERU on the CIFAR100 dataset. Additionally, HMID-Net achieves an accuracy of 28.6% on the ImageNet dataset, surpassing CLIP and MERU by 5.1% and 4.3%, respectively. On the MNIST and SST2 datasets, HMID-Net achieves performance on par with that of CLIP and MERU. As noted by (Desai et al. 2023), HMID-Net exhibits relatively suboptimal performance on datasets with fewer covered concepts, such as SST2, which is derived from movie reviews. Pretraining on larger datasets may improve performance. Overall, HMID-Net is highly competitive compared to Euclidean space-based methods.
@@
Image and text retrieval
Ablations experiments
Masking ratio. We first conducted a study on image masking ratios based solely on MERU, as shown in Table LABEL:mask. The 0% masking ratio refers to our baseline MERU. A 50% masking ratio yields optimal performance, with a 1.7% improvement on ImageNet and a 1.3% improvement on COCO for the T2I retrieval task. Compared to BERT’s 15% masking ratio, images have significant pixel redundancy, enabling a higher masking ratio. However, a 75% masking ratio results in a performance decline due to the loss of critical information, which hinders contrastive learning. This effect is similar to the findings in FLIP (Li et al. 2023). Unless specified otherwise, we use a default masking ratio of 50%.
@@
Qualitative analysis
This section presents a qualitative analysis of the visual-semantic hierarchy. General objects are closer to the [Root], while specific objects are near the boundary (Ramasinghe et al. 2024). The distance from the origin indicates uncertainty, useful for retrieval tasks. In this hierarchy, text is closer to the origin and images nearer the boundary. For example, in Fig. 1(a), the ”Cheerful smile” is near the [ROOT], while the image is positioned closer to the boundary.
Our experiment closely follows MERU (Desai et al. 2023). A subset of images is randomly selected from Pixels, with textual descriptions retrieved from a curated set of 750 captions on pexels.com. We interpolate 50 steps along the geodesic between the image embedding and the [ROOT], selecting the textual description with the highest Lorentzian inner product at each step. Duplicates are removed, and the top five descriptions are retained. Fig. 6 shows that both HMID-Net and MERU effectively capture the visual-semantic hierarchy, with descriptions becoming more generic closer to the [ROOT].
Discussion and Conclusion
In this paper, we propose the Hyperbolic Masked Image and Distillation Network (HMID-Net), which integrates Masked Image Modeling (MIM) and knowledge distillation in hyperbolic space to more efficiently learn the visual-semantic hierarchy. Our method introduces knowledge distillation to hyperbolic space, achieving training efficiency on par with Euclidean space. Experimental results demonstrate that HMID-Net enhances real-world understanding and outperforms baseline models MERU and CLIP across a range of tasks, highlighting the effectiveness of MIM and knowledge distillation in hyperbolic space. However, our method still have several limitations. First, HMID-Net exhibits slower convergence during training, which we attribute to the difficulty of jointly optimizing multiple loss functions. The presence of multiple objectives increases the complexity of the optimization landscape, making it challenging for the model to locate a single optimal solution that balances all tasks effectively. Second, the performance of HMID-Net on datasets such as SST-2 remains suboptimal. We believe this is primarily due to the limited coverage of SST-2-like samples in the current training dataset, which hinders the model’s ability to generalize to sentiment classification tasks that require sensitivity to fine-grained emotional and linguistic cues.
In future work, to address the slow convergence issue, we plan to refine the knowledge distillation loss function, aiming to reduce conflicts among multiple objectives and facilitate more stable and efficient optimization. By better aligning the learning signals from different supervision sources, we expect to accelerate convergence during training. To improve the generalization performance on datasets such as SST-2, we intend to expand and diversify the training data by incorporating large-scale, sentiment-rich datasets. This enhancement will allow the model to learn from a broader distribution of linguistic patterns and emotional expressions, thereby improving its adaptability and performance across a wider range of downstream tasks.